OpenStudy (anonymous):

for f(x)=x+3 and g(x)=x^2 find (g o f)(1)

OpenStudy (anonymous):

The answer is 16 (g o f)(1) = g( f(1)) = g ( 4) = 16

OpenStudy (anonymous):

sure?

OpenStudy (jkristia):

yes I get that too (g o f)(x) = (x+3)^2 = (1+3)^4 = 16